Sales forecasts. For the prior three​ years, sales for California Cement Company have been ​$20,068,000 ​(2011), $21,073,000 ​(2012), and $22,809,000 ​(2013). The company uses the prior two​ years' average growth rate to predict the coming​ year's sales. What were the sales growth rates for 2012 and​ 2013? What is the expected sales growth rate using a​ two-year average for​ 2014? What is the sales forecast for​ 2014?

What was the sales growth rate for​ 2012?

​(Round to two decimal​ places.)

What was the sales growth rate for​ 2013?

​(Round to two decimal​ places.)

What is the expected sales growth rate using a​ two-year average for​ 2014?

​(Round to two decimal​ places.)

What is the sales forecast for​ 2014?

​ (Round to the nearest​ dollar.)
Answer #1

Answerr:

Given

2011 sales S1=$20068000

2012 sales S2=$21073000

2013 sales S3=$22809000

A)

Sales growth rate for 2012 G1= (S2-S1)/S1=(21073000-20068000)/20068000=5.01%

B)

Sales growth rate for 2013 G2= (S3-S2)/S2=(22809000-21073000)/21073000=8.24%

C)

Expected growth rate for 2014 R= (G1+G2)/2=(5.01%+8.24%)/2=6.62%

D)

Sales forecast for 2014 =S3*(1+R)=22809000*(1+6.62%)=$24319640.4